OpenStudy (anonymous):

Identify the rule for the function table. input x 24 12 -12 output y 12 6 -6 possible answers: A. Y=12-X B. Y=X-12 C. Y=X / 2 D. Y = 2 / X WILL MEDAL! Thanks

OpenStudy (anonymous):

PLEASE HELP!

Mehek (mehek14):

(24, 12) (12, 6) \(\dfrac{6 - 12}{12 - 24} = \dfrac{1}{2}\)

OpenStudy (anonymous):

sorry i dont understand.. this is confusing to me

Mehek (mehek14):

so it would be C because \(\dfrac{24}{2} = 12 \\ \dfrac{12}{2} = 6 \\ \dfrac{-12}{2} = -6\)

Mehek (mehek14):

you divide the value of x by 2 and get y that would be the function for your table

Mehek (mehek14):

you saw for the x value you had 24 you divided that by 2 and got the y value , 12 same goes for the res

OpenStudy (anonymous):

i am trying.. I also need to complete a function table but no idea how :( input x 32 14 ? -2 -10 output y 20 2 -6 -14 ? it is asking to show the function rule.. can you help me? thanks

Mehek (mehek14):

got options?

OpenStudy (anonymous):

no

Mehek (mehek14):

actually if you look at it, all you're doing is subtracting by 12

Mehek (mehek14):

so to find the first ? add 12 + (-6) = 6 second ? -10 - 12 = -22

OpenStudy (anonymous):

yes i got that also but what does it mean by show a function rule?

Mehek (mehek14):

y = x - 12 you subtract 12 from the x value

OpenStudy (anonymous):

ahhh i see.. thank you so much!